factor completely
25m^2+81-90m

Answer by jim_thompson5910(13679)

Start with the given expression.


Rearrange the terms in descending exponential order.



Looking at the expression , we can see that the first coefficient is , the second coefficient is , and the last term is .


Now multiply the first coefficient by the last term to get .


Now the question is: what two whole numbers multiply to (the previous product) and add to the second coefficient ?


To find these two numbers, we need to list all of the factors of (the previous product).


Factors of :
1,3,5,9,15,25,27,45,75,81,135,225,405,675,2025
-1,-3,-5,-9,-15,-25,-27,-45,-75,-81,-135,-225,-405,-675,-2025


Note: list the negative of each factor. This will allow us to find all possible combinations.


These factors pair up and multiply to .
1*2025
3*675
5*405
9*225
15*135
25*81
27*75
45*45
(-1)*(-2025)
(-3)*(-675)
(-5)*(-405)
(-9)*(-225)
(-15)*(-135)
(-25)*(-81)
(-27)*(-75)
(-45)*(-45)

Now let's add up each pair of factors to see if one pair adds to the middle coefficient :


First NumberSecond NumberSum
120251+2025=2026
36753+675=678
54055+405=410
92259+225=234
1513515+135=150
258125+81=106
277527+75=102
454545+45=90
-1-2025-1+(-2025)=-2026
-3-675-3+(-675)=-678
-5-405-5+(-405)=-410
-9-225-9+(-225)=-234
-15-135-15+(-135)=-150
-25-81-25+(-81)=-106
-27-75-27+(-75)=-102
-45-45-45+(-45)=-90



From the table, we can see that the two numbers and add to (the middle coefficient).


So the two numbers and both multiply to and add to


Now replace the middle term with . Remember, and add to . So this shows us that .


Replace the second term with .


Group the terms into two pairs.


Factor out the GCF from the first group.


Factor out from the second group. The goal of this step is to make the terms in the second parenthesis equal to the terms in the first parenthesis.


Combine like terms. Or factor out the common term


Condense

---------------------------------------------


Answer:


So factors to .


In other words,


Note: you can check the answer by FOILing to get or by graphing the original expression and the answer (the two graphs should be identical).